OpenStudy (anonymous):

sand is falling into a cone whose height is always 4 times the diameter. if the sand is falling at a rate of 500 cubic meters per hours how fast is the radius increasing when the volume is 1000pi/3 cubic meters

OpenStudy (anonymous):

Let r = radius , V = volume, h = height h = 8r V =1/12( pi.r^2.h) V = 1/12(pi.r^2.8r) V= 2/3(pi.r^3) dv/dr = 4/3.pi.r^2 it is given that dv/dr is 1000pi/3 1000pi/3 =4/3.pi.r^2 1000 = 4r^2 r = square root of 250 meters the radius is increasing at a rate of square root of 250 meters per hour